# Impurity dynamics in a zero-temperature gas

### v1

@@ Abstract
If energy is suddenly released in a localized region of space uniformly filled with identical stationary hard spheres of radii $a$, the outcome is a blast with an asymptotically spherical shock wave separating moving and stationary hard spheres. The radius of the region filled with the moving spheres grows as $R\sim(Et^{2}/n_{\infty})^{1/(d+2)}$, where $d$ is the spatial dimension, $n_{\infty}$ the number density of stationary spheres and $E$ is the total energy injected. The simplest way to inject energy is to kick a few ‘impurity’ particles. Using hydrodynamics and kinetic theory, we argue that the typical displacement of an impurity scales as $R_{\rm imp}\sim\lambda(R/\lambda)^{(4+3d^{2})/(8+3d^{2})}$, where $\lambda\sim 1/(n_{\infty}a^{d-1})$ is the mean-free path in the initial state. The number of collisions experienced by each impurity is $(R/\lambda)^{(8+2d^{2})/(8+3d^{2})}$, while its average speed decreases as $t^{-d(8-2d+3d^{2})/[(2+d)(8+3d^{2})]}$. In $2D$, the predictions for impurity displacement, collision numbers, and speed are $t^{2/5},~{}t^{2/5}$ and $t^{-2/5}$, respectively. These predictions are in reasonable agreement with the results of molecular dynamics simulations.
@@ I Introduction
A localized instantaneous release of energy generates a blast, first investigated in the context of atomic explosions Taylor (1950a, b); Sedov (1959); von Neumann (1963). The blast is strong as long as the pressure behind the shock wave greatly exceeds the atmospheric pressure. For the strong blast, Euler equations describing an ideal compressible gas admits an exact solution; see textbooks Zeldovich and Raiser (1967); Landau and Lifshitz (1987); Barenblatt (1996). The shock wave eventually weakens, and the problem becomes analytically intractable. However, we look at the idealized case where the pressure in front of the shock is zero, so the blast is infinitely strong throughout the evolution. Infinitely strong blasts mimic supernova explosions and other astrophysical phenomena Ostriker and McKee (1988); Tang and Chevalier (2017); Vink (2020); Modelevsky and Sari (2021).
@@
One can examine the infinitely strong blast problem at a microscopic level by considering a hard sphere gas with particles initially at rest, i.e., at zero temperature. How does this system evolve when energy is injected into a localized region, for example, by giving energy to a few particles? An infinite system with initially immobile particles has an additional virtue—it is amenable to direct molecular dynamics simulations since there is only a finite number of moving particles. Hence, one can investigate the behavior in a numerically precise way. This system thus provides a fruitful laboratory for probing the validity of hydrodynamics, explaining the popularity of the setting with initially immobile hard spheres and the number of recent papers studying the blast for this system Antal et al. (2008); Barbier et al. (2015, 2016); Joy et al. (2021); Joy and Rajesh (2021); Kumar and Rajesh (2022); Chakraborti et al. (2021); Ganapa et al. (2021); Chakraborti et al. (2022); Kumar and Rajesh (2024, 2025a); Singh et al. (2023). These studies compared the predictions of Euler hydrodynamics with the results from microscopic simulations. Somewhat surprisingly, one finds rather good agreement. However, deviations from the predictions of the Euler equations have been observed. These deviations can be understood by taking into account dissipative effects such as heat conduction in a core region around the center of the blast, i.e., one needs to use the full Navier-Stokes equations Chakraborti et al. (2021); Ganapa et al. (2021); Singh et al. (2023).
@@
Previous studies focused on coarse-grained observables such as density, temperature, and pressure. The emergence of the hydrodynamic behavior is questionable, at least at first sight, as the gas outside the shock wave has zero temperature. The success of hydrodynamics comes as a pleasant surprise. Going beyond the scope of conventional hydrodynamics predicting macroscopic characteristics, one can study mesoscopic characteristics such as the self-diffusion phenomenon inside a blast. The present work explores precisely this question. Employing hydrodynamics and kinetic theory arguments we deduce scaling laws describing the characteristics of the impurity.
@@ II The model and summary of results
Consider a homogeneous hard-sphere gas with initially static spheres, each of radius $a$, and unit mass $m=1$. We perturb the system at time $t=0$ by giving energy, $E$, to a few particles Antal et al. (2008); Krapivsky et al. (2010). We will refer to these particles as the impurities. The impurities elastically collide with the host spheres, which begin to move and elastically collide with other spheres, generating a growing number of moving spheres. We are primarily interested in the ‘fate’ of the impurities, specifically the asymptotic distributions of their position and velocity. In Fig. 1, we present a simulation of a blast in a two-dimensional hard-disc gas initiated by injecting energy into four impurity particles. The figure also shows the trajectories of these impurity particles against the backdrop of the evolving blast environment.
@@
where $n_{\infty}$ is the number density of spheres and $d$ is the spatial dimension. At first sight, the sub-linear growth $R\sim t^{2/(d+2)}$ is surprising because the shock wave always moves faster than the sonic speed Landau and Lifshitz (1987). However, the sonic speed vanishes in the zero-temperature gas, so, albeit the speed of the shock decays as $t^{-d/(d+2)}$, the shock remains infinitely strong throughout the evolution. The position of the shock wave depends only on macroscopic variables, i.e., total energy, $E$, and number density, $n_{\infty}$. The notation $A_{d}$ for the amplitude in Eq. (1) emphasizes the dependence on the spatial dimension, but $A_{d}$ also depends on the volume fraction
@@
occupied by spheres. These amplitudes are generally unknown since even the equation of state for the dense hard-sphere gas is unknown. In the $\phi\to 0$ limit, the hard-sphere gas is dilute and ideal. Euler equations for ideal compressible gas behind the shock arising in our infinitely strong blast problem admit an exact solution for arbitrary adiabatic index $\gamma$. We set $\gamma=1+2/d$, the adiabatic index characterizing ideal monoatomic gases Resibois and Leener (1977); Dorfman et al. (2021); Lifshitz and Pitaevskii (1981). In the physically relevant dimensions, the amplitudes are
[…]
However, the hard-sphere simulations revealed deviations from the exact solution of the Euler equations for the ideal compressible gas. One should account for heat conduction in the core region which grows slower than the shock wave, i.e., rely on the Navier-Stokes equations Chakraborti et al. (2021); Ganapa et al. (2021); Singh et al. (2023). There is no sharp boundary separating the core from the bulk described by Euler equations, but one can still talk about the characteristic radius of the core region. The temperature in the core decays with time, yet the ratio of this temperature to the temperature in the bulk of the moving gas behind the shock diverges with time. Thus, the core region is ‘hot’; we denote its characteristic radius by $H$. This radius grows as Singh et al. (2023)
$H=\lambda\left(\frac{R}{\lambda}\right)^{h_{d}},$ (3)
and we revisit the derivation of this formula in the next section.
@@
    $R_{\text{imp}}=\lambda\left(\frac{R}{\lambda}\right)^{h_{d}},\qquad h_{d}=\frac{4+3d^{2}}{8+3d^{2}},$ (4)
    where the mean-free path is given by
    $\lambda=\frac{1}{n_{\infty}\Omega_{d-1}a^{d-1}}.$ (5)
    In two dimensions, this yields $R_{\text{imp}}\propto t^{2/5}$. We test this prediction in simulations and a comparison with the theory is shown in Fig. 2(A).
@@
    $V_{\text{imp}}=V\left(\frac{a}{\lambda}\right)^{\frac{d-1}{2}}\left(\frac{\lambda}{R}\right)^{\omega_{d}},\quad\omega_{d}=\frac{d}{2}-\frac{d^{2}}{8+3d^{2}}.$ (6)
@@
    $V_{\text{imp}}=V\phi^{\frac{d-1}{2}}\left(\frac{\lambda}{R}\right)^{\omega_{d}}.$ (7)
@@
    $u_{\text{imp}}=V\phi^{\frac{d-1}{2}}\left(\frac{\lambda}{R}\right)^{\frac{d}{2}+1-h_{d}}.$ (8)
    For two dimensions, this gives $u_{\text{imp}}\propto t^{-3/5}$. A comparison between theory and simulation is shown in Fig. 2(D).
@@
  • As part of our heurisitic derivation of some of the above results, we also show that in the core region, the density $n_{0}$ and temperature $T_{0}$ are given by
@@
    $T_{0}=mV^{2}\left(\frac{a}{\lambda}\right)^{d-1}\left(\frac{\lambda}{R}\right)^{2\omega_{d}}.$ (11)
    In two dimensions, $n_{0}\propto t^{-1/5}$ and $T_{0}\propto t^{-4/5}$; see Fig. 2(E)-(F) for the comparison with simulation results.
@@
The mean free path depends on the radius of spheres, a microscopic variable. Therefore, in contrast to the position of the shock wave, (1), predicted by ideal compressible hydrodynamics, the determination of the displacement of the impurity seemingly requires a microscopic framework. Indeed, the derivation of (4) relies on the arguments used in kinetic theory Resibois and Leener (1977); Dorfman et al. (2021). Remarkably, the length scale $R_{\text{imp}}$ is identical to the one mentioned in (3), that arises in a more accurate hydrodynamic description of the flow in the core of the blast. Thus, the impurity remains inside the core.
@@
The above formulas apply when $d\geq 2$, particularly in two and three dimensions. The one-dimensional setting is pathological due to the lack of hydrodynamic behavior: There is no mixing since two elastically colliding particles exchange the velocities and effectively pass through each other. In our setting, there is only one moving particle at any moment. One circumvents this pathology by allowing for distributed particle masses. The hydrodynamic behavior then emerges Antal et al. (2008); Chakraborti et al. (2021); Ganapa et al. (2021). The core region was originally detected in one dimension Chakraborti et al. (2021); Ganapa et al. (2021). The behavior of the impurity remains pathological since, in one dimension, the impurity is forever caged between the right and left neighbors. This caging property can be circumvented by postulating that the impurity colliding with a host particle passes with a certain probability $p$, while elastic collision occurs with probability $1-p$. (This trick allows one to devise a meaningful version of the one-dimensional Lorentz gas Piasecki and Wajnryb (1979); Krapivsky and Redner (1997).) With such an amendment, the displacement of the impurity should be comparable with the size of the core region. Below, we always assume that $d\geq 2$ and employ the basic setting with spheres of equal masses undergoing elastic collisions.
@@
Figure 2: Time evolution of various impurity and fluid properties plotted on log-log scale: (A) impurity displacement $R_{\text{imp}}$; (B) impurity speed $V_{\text{imp}}$; (C) impurity collision count $C_{\text{imp}}$; (D) fluid speed around the impurity $u_{\text{imp}}$; (E) central density $n_{0}$; (F) central temperature $T_{0}$. All quantities are asymptotically expected to have power-law time dependence of the form $t^{\alpha}$ and the insets show the running time slopes plotted with time. The dark dashed lines correspond to the theoretical predictions.
@@ III Heuristic Derivations
The strategy of our heuristic derivation is that we first establish the properties in the core region where dissipation (viscosity and thermal conductivity) is important. In particular, we find the typical size (3) of the core region, and the density (10) and temperature (11) in this region. Next, to justify the scaling laws (4) and (7) –(9), we assume that in the long time limit, the impurity is in the core region and demonstrate self-consistency thus confirming a posteriori that the impurity is inside the core. The derivation of Eqs. (4) and (7) –(9) then follows from arguments based on elementary kinetic theory.
The growth law $H\propto R^{h_{d}}$ for the characteristic size of the core region was established in Singh et al. (2023). We now derive the more accurate result (3). (Below, we also use the same approach to generalize (3) to monatomic gases with point particles and algebraically diverging repulsive interactions.) To estimate the size of the core region where dissipative effects matter, it suffices to keep heat conduction and ignore viscous dissipation Singh et al. (2023); Chakraborti et al. (2021); Ganapa et al. (2021). In this situation, the entropy equation reads
@@
where we ignore (analytically unknown) numerical factor. Using (13) we recast (12) into
$\rho^{1+2/d}\,(\partial_{t}+v\partial_{r})\frac{T}{\rho^{2/d}}=\frac{1}{r^{d-1}}\,\partial_{r}\big{(}r^{d-1}\kappa\,\partial_{r}T\big{)}$ (15)
Inserting (14) into (15) we estimate the radius $H$ of the core where the heat transfer is comparable with the terms on the left-hand side
$\rho\sim n_{\infty}\lambda\,\sqrt{T}\,\frac{t}{H^{2}}$ (16)
From the exact solution outside the core Singh et al. (2023) one finds the asymptotic behaviors valid when $r\ll R$:
$\frac{\rho}{n_{\infty}}$ $=\left(\frac{r}{R}\right)^{\frac{d^{2}}{2}}$ (17a)
$T$ $=\left(\frac{R}{t}\right)^{2}\left(\frac{r}{R}\right)^{-\frac{d^{2}}{2}}$ (17b)
The divergence of the temperature when $r\to 0$ is physically impossible. The asymptotic behaviors (17a)–(17b) are valid for $r>H$. Setting $r=H$ in (17a)–(17b) and inserting into (16) we arrive at (3).
@@
$T_{0}=V^{2}\left(\frac{a}{\lambda}\right)^{d-1}\,\left(\frac{\lambda}{R}\right)^{2\omega_{d}}$ (19)
@@
$D_{0}=\lambda\,\frac{n_{\infty}}{n_{0}}\,V_{\text{imp}}=\lambda V\phi^{\frac{d-1}{2}}\,\left(\frac{\lambda}{R}\right)^{3\omega_{d}-d}$ (20)
The typical displacement of the impurity is found from $R_{\text{imp}}=\sqrt{D_{0}t}$ which in conjunction with Eqs. (20), (2), and (1) lead to the announced estimate (4).
Thus assuming that the impurity stays within the core region we have demonstrated the consistency by showing that the typical displacement of the impurity grows as the size of the core region. The impurity certainly stays behind the shock wave, but assuming that it explores the $r\leq R(t)$ ball and repeating the above line of reasoning gives $R_{\text{imp}}$ that grows slower than $R\propto t^{2/(d+2)}$. This failure has led us to the $R_{\text{imp}}\sim H$ guess that turned out to be self-consistent.
@@
$u_{\text{imp}}=V\phi^{\frac{d-1}{2}}\left(\frac{\lambda}{R}\right)^{\frac{d}{2}+1-h_{d}}$ (21)

### v2

@@ Abstract
If energy is suddenly released in a localized region of space uniformly filled with identical stationary hard spheres, the outcome is a blast with an asymptotically spherical shock wave separating moving and stationary hard spheres. The radius $R(t)$ of the region filled with the moving spheres grows as $t^{2/(d+2)}$, where $d$ is the spatial dimension. The simplest way to inject energy is to kick a few ‘impurity’ particles. Using hydrodynamics and kinetic theory, we argue that the typical displacement of an impurity scales as $R_{\rm imp}\sim\lambda(R/\lambda)^{(4+3d^{2})/(8+3d^{2})}$, where $\lambda$ is the mean-free path in the initial state. The number of collisions experienced by each impurity grows as $(R/\lambda)^{(8+2d^{2})/(8+3d^{2})}$, while its average speed decreases as $t^{-d(8-2d+3d^{2})/[(2+d)(8+3d^{2})]}$. In $2D$, the predictions for impurity displacement, collision numbers, and speed are $t^{2/5},~t^{2/5}$ and $t^{-2/5}$, respectively. These predictions are in reasonable agreement with the results of molecular dynamics simulations.
@@ I Introduction
A localized instantaneous release of energy generates a blast, first investigated in the context of atomic explosions Taylor (1950a, b); Sedov (1959); von Neumann (1963). The blast is strong as long as the pressure behind the shock wave greatly exceeds the atmospheric pressure. For the strong blast, Euler equations describing an ideal compressible gas admit an exact solution; see textbooks Zeldovich and Raiser (1967); Landau and Lifshitz (1987); Barenblatt (1996). The shock wave eventually weakens, and the problem becomes analytically intractable. However, we look at the idealized case where the pressure in front of the shock is zero, so the blast is infinitely strong throughout the evolution. Infinitely strong blasts mimic supernova explosions and other astrophysical phenomena Ostriker and McKee (1988); Tang and Chevalier (2017); Vink (2020); Modelevsky and Sari (2021).
@@
One can examine the infinitely strong blast problem at a microscopic level by considering a hard sphere gas with particles initially at rest, i.e., at zero temperature. How does this system evolve when energy is injected into a localized region, for example, by giving energy to a few particles? An infinite system with initially immobile particles has an additional virtue—it is amenable to direct molecular dynamics simulations since there is only a finite number of moving particles. Hence, one can investigate the behavior in a numerically precise way. This system thus provides a fruitful laboratory for probing the validity of hydrodynamics, explaining the popularity of the setting with initially immobile hard spheres and the number of recent papers studying the blast for this system Antal et al. (2008); Barbier et al. (2015, 2016); Joy et al. (2021); Joy and Rajesh (2021); Kumar and Rajesh (2022); Chakraborti et al. (2021); Ganapa et al. (2021); Chakraborti et al. (2022); Kumar and Rajesh (2025b, a); Singh et al. (2023). These studies compared the predictions of Euler hydrodynamics with the results from microscopic simulations. Somewhat surprisingly, one finds rather good agreement. However, deviations from the predictions of the Euler equations have been observed. These deviations can be understood by taking dissipative effects into account, such as heat conduction in a core region around the center of the blast; i.e., one needs to use the full Navier-Stokes equations Chakraborti et al. (2021); Ganapa et al. (2021); Singh et al. (2023).
@@
Previous studies focused on coarse-grained observables such as density, temperature, and pressure. The emergence of the hydrodynamic behavior is questionable, at least at first sight, as the gas outside the shock wave has zero temperature. The success of hydrodynamics comes as a pleasant surprise. Going beyond the scope of conventional hydrodynamics predicting macroscopic characteristics, one can study mesoscopic characteristics such as the self-diffusion phenomenon inside a blast. The present work explores precisely this question. Employing hydrodynamics and kinetic theory arguments, we deduce scaling laws describing the characteristics of the impurity.
@@ II The model and summary of results
Consider a homogeneous gas of hard spheres that are initially at rest, with number density $n_{\infty}$ and mass density $\rho_{\infty}$. Each hard sphere has radius $a$ and unit mass ($m=1$); therefore, the mass density and number density are related by $\rho=n$. We perturb the system at time $t=0$ by giving total energy, $E$, to a few particles Antal et al. (2008); Krapivsky et al. (2010). We will refer to these particles as the impurities. The impurities elastically collide with the host spheres, which begin to move and elastically collide with other spheres, generating a growing number of moving spheres. We are primarily interested in the ‘fate’ of the impurities, specifically the asymptotic distributions of their position and velocity. In Fig. 1, we present a simulation of a blast in a two-dimensional hard-disc gas initiated by injecting energy into four impurity particles. The figure also shows the trajectories of these impurity particles against the backdrop of the evolving blast environment.
@@
where $n_{\infty}$ is the number density (a mass prefactor $m$ has been set to one) of the quiescent spheres and $d$ is the spatial dimension. At first sight, the sub-linear growth $R\sim t^{2/(d+2)}$ is surprising because the shock wave always moves faster than the sonic speed Landau and Lifshitz (1987). However, the sonic speed vanishes in the zero-temperature gas, so, albeit the speed of the shock decays as $t^{-d/(d+2)}$, the shock remains infinitely strong throughout the evolution. The position of the shock wave depends only on macroscopic variables, i.e., total energy, $E$, and number density, $n_{\infty}$. The notation $A_{d}$ for the amplitude in Eq. (1) emphasizes the dependence on the spatial dimension, but $A_{d}$ also depends on the volume fraction
@@
occupied by spheres. The analytical expression for amplitudes are generally unknown since even the equation of state for the dense hard-sphere gas is not known in closed form. However, they can readily be computed numerically. In the $\phi\to 0$ limit, the hard-sphere gas is dilute and ideal. Euler equations for ideal compressible gas behind the shock arising in our infinitely strong blast problem admit an exact solution for arbitrary adiabatic index $\gamma$. We set $\gamma=1+2/d$, the adiabatic index characterizing ideal monoatomic gases Resibois and Leener (1977); Dorfman et al. (2021); Lifshitz and Pitaevskii (1981). In the physically relevant dimensions, the amplitudes are
$A_{2}=0.351\,935\,906\ldots,\quad A_{3}=0.072\,401\,066\ldots$ (3)
The work of TvNS showed that the Euler equations, for the hydrodynamic fields for density ($\rho$), velocity ($u$), and temperature ($T$), admitted self-similar scaling forms in terms of the scaling variable $r/R(t)$. For the ideal gas, the scaling functions can be determined analytically in all dimensions Singh et al. (2023)—we will refer to these as the TvNS solution. However, hard-sphere simulations in Refns. Chakraborti et al. (2021); Ganapa et al. (2021); Singh et al. (2023) have revealed deviations from the solution of the Euler equations for the compressible gas. They showed that the TvNS predicted growth law in Eq. (1) was accurate and the scaling solutions were valid in a large volume of the gas behind the shock front, which we will refer to as the bulk. However, it was found that in a smaller region centered around the blast center, one needs to account for heat conduction. In this core region, which grows slower than the shock wave, one has to rely on the Navier-Stokes equations Chakraborti et al. (2021); Ganapa et al. (2021); Singh et al. (2023). While there is no sharp boundary separating the core from the bulk described by Euler equations, one can still talk about the characteristic radius of the core region. The temperature in the core decays with time, yet the ratio of this temperature to the temperature in the bulk of the moving gas behind the shock diverges with time. Thus, the core region is ‘hot’; we denote its characteristic radius by $H$. This radius grows as Singh et al. (2023)
$H=\lambda\left(\frac{R}{\lambda}\right)^{h_{d}},\qquad h_{d}=\frac{4+3d^{2}}{8+3d^{2}},$ (4)
and $\lambda$ is the mean-free path for the quiescent hard sphere gas, given by
$\lambda=\frac{1}{n_{\infty}\Omega_{d-1}a^{d-1}}.$ (5)
We revisit the derivation of this formula in the next section.
@@
    $R_{\text{imp}}=\lambda\left(\frac{R}{\lambda}\right)^{h_{d}}.$ (6)
    In two-dimensions, this yields $R_{\text{imp}}\propto t^{2/5}$. We test this prediction in simulations, and a comparison with the theory is shown in Fig. 2(A).
@@
    $V_{\text{imp}}=\sqrt{E}\,\left(\frac{a}{\lambda}\right)^{\frac{d-1}{2}}\left(\frac{\lambda}{R}\right)^{\omega_{d}},\quad\omega_{d}=\frac{d}{2}-\frac{d^{2}}{8+3d^{2}}.$ (7)
@@
    $V_{\text{imp}}=\sqrt{E}\,\phi^{\frac{d-1}{2}}\left(\frac{\lambda}{R}\right)^{\omega_{d}}.$ (8)
@@
    $u_{\text{imp}}=\sqrt{E}\phi^{\frac{d-1}{2}}\left(\frac{\lambda}{R}\right)^{\frac{d}{2}+1-h_{d}}.$ (9)
    For two-dimensions, this gives $u_{\text{imp}}\propto t^{-3/5}$. A comparison between theory and simulation is shown in Fig. 2(D).
@@
  • As part of our heuristic derivation of some of the above results, we also show that in the core region, the characteristic density $n_{0}$ and temperature $T_{0}$ are given by
@@
    $T_{0}=E\left(\frac{a}{\lambda}\right)^{d-1}\left(\frac{\lambda}{R}\right)^{2\omega_{d}}.$ (12)
    In two-dimensions, $n_{0}\propto t^{-1/5}$ and $T_{0}\propto t^{-4/5}$; see Fig. 2(E)-(F) for the comparison with simulation results.
@@
The mean free path depends on the radius of spheres, a microscopic variable. Therefore, in contrast to the position of the shock wave, (1), predicted by ideal compressible hydrodynamics, the determination of the displacement of the impurity seemingly requires a microscopic framework. Indeed, the derivation of (6) relies on the arguments used in kinetic theory Resibois and Leener (1977); Dorfman et al. (2021). Remarkably, the length scale $R_{\text{imp}}$ is identical to the one mentioned in (4), that arises in a more accurate hydrodynamic description of the flow in the core of the blast. Thus, the impurity remains inside the core. Note that the core-bulk boundary is not sharp, and what the previous line implies is that the impurity displacement and the core size have the same scaling with time.
@@
The above formulas apply when $d\geq 2$, particularly in two and three dimensions. The one-dimensional setting is pathological due to the lack of hydrodynamic behavior: There is no mixing since two elastically colliding particles exchange the velocities and effectively pass through each other. In our setting, there is only one moving particle at any moment. One circumvents this pathology by allowing for distributed particle masses. The hydrodynamic behavior then emerges Antal et al. (2008); Chakraborti et al. (2021); Ganapa et al. (2021). The core region was originally detected in one dimension Chakraborti et al. (2021); Ganapa et al. (2021). The behavior of the impurity remains pathological since, in one dimension, the impurity is forever caged between the right and left neighbors. This caging property can be circumvented by postulating that the impurity colliding with a host particle passes with a certain probability $p$, while elastic collision occurs with probability 1p1-p. This trick allows one to devise a meaningful version of the one-dimensional Lorentz gas Piasecki and Wajnryb (1979); Krapivsky and Redner (1997); van Beijeren (1982). With such an amendment, the displacement of the impurity should be comparable with the size of the core region. Alternatively, one can consider a quasi-1D geometry where particles are allowed to overtake each other. Below, we always assume that $d\geq 2$ and employ the basic setting with spheres of equal masses undergoing elastic collisions.
@@
Figure 2: Time evolution of various impurity and fluid properties plotted on log-log scale in case of a 22-dimensional hard-sphere gas: (A) impurity displacement $R_{\text{imp}}$; (B) impurity speed $V_{\text{imp}}$; (C) impurity collision count $C_{\text{imp}}$; (D) fluid speed around the impurity $u_{\text{imp}}$; (E) central density $n_{0}$; (F) central temperature $T_{0}$. All quantities are asymptotically expected to have power-law time dependence of the form $t^{\alpha}$ and the insets show the running temporal exponents plotted with time (see Sec. (V) for details). The dark dashed lines correspond to the theoretical predictions. All figures and their insets include error bars. However, the estimated statistical uncertainties are smaller than the symbol size and the line thickness used in the plots; as a result, the error bars are not visually discernible.
@@ III Heuristic Derivations
The strategy of our heuristic derivation is that we first establish the properties in the core region where dissipation (viscosity and thermal conductivity) is important. Since we are primarily interested in scaling laws, throughout the derivation we will typically ignore many pre-factors. In particular, we find the typical size (4) of the core region, and the density (11) and temperature (12) in this region. Next, to justify the scaling laws (6) and (8)–(10), we assume that in the long time limit, the impurity is in the core region and demonstrate self-consistency, thus confirming a posteriori that the impurity is inside the core. The derivation of Eqs. (6) and (8)–(10) then follows from arguments based on elementary kinetic theory. The growth law $H\propto R^{h_{d}}$ for the characteristic size of the core region was established in Singh et al. (2023). We now derive the more accurate result (4).
To estimate the size of the core region where dissipative effects matter, it suffices to keep heat conduction and ignore viscous dissipation Singh et al. (2023); Chakraborti et al. (2021); Ganapa et al. (2021). In this situation, the entropy equation reads
@@
where we ignore the numerical pre-factor. Using (14) we recast (13) into
$\rho^{1+2/d}\,(\partial_{t}+u\partial_{r})\frac{T}{\rho^{2/d}}=\frac{1}{r^{d-1}}\,\partial_{r}\big(r^{d-1}\kappa\,\partial_{r}T\big)$ (16)
To estimate the radius $H$ of the core, defined as the region where the heat transfer is comparable with the terms on the left-hand side, we first insert (15) into (16). Approximating $\partial_{t}T\sim T/t,\partial_{r}T\sim T/H$ and $u\sim H/t$ we get the following relation between the size of the core $H$, time $t$, and the core density, $\rho$:
$\rho\sim n_{\infty}\lambda\,\sqrt{T}\,\frac{t}{H^{2}}.$ (17)
To find estimates of $\rho$ and $T$, we use the TvNS solution which is expected to be valid for $r\gtrsim H$. From the exact solution outside the core Singh et al. (2023) one finds the asymptotic behaviors valid when $H\lesssim r\ll R$:
$\frac{\rho}{n_{\infty}}$ $\sim\left(\frac{r}{R}\right)^{\frac{d^{2}}{2}}$ (18a)
$T$ $\sim\left(\frac{R}{t}\right)^{2}\left(\frac{r}{R}\right)^{-\frac{d^{2}}{2}}$ (18b)
The divergence of the temperature when $r\to 0$ is physically impossible. The asymptotic behaviors (18a)–(18b) are valid for $r\gtrsim H$. Setting $r=H$ in (18a)–(18b) and inserting into (17) we arrive at (4).
@@
$T_{0}=E\,\left(\frac{a}{\lambda}\right)^{d-1}\,\left(\frac{\lambda}{R}\right)^{2\omega_{d}}$ (20)
@@
$D_{0}=\lambda\,\frac{n_{\infty}}{n_{0}}\,V_{\text{imp}}=\lambda\sqrt{E}\,\phi^{\frac{d-1}{2}}\,\left(\frac{\lambda}{R}\right)^{3\omega_{d}-d}$ (21)
The typical displacement of the impurity is found from $R_{\text{imp}}\sim\sqrt{D_{0}t}$ which in conjunction with Eqs. (21), (2), and (1) lead to the announced estimate (6).
Thus assuming that the impurity stays within the core region, we have demonstrated the consistency by showing that the typical displacement of the impurity grows as the size of the core region. The impurity certainly stays behind the shock wave, but assuming that it explores the $r\leq R(t)$ ball and repeating the above line of reasoning gives $R_{\text{imp}}$ that grows slower than $R\propto t^{2/(d+2)}$. This failure has led us to the $R_{\text{imp}}\sim H$ guess that turned out to be self-consistent.
@@
$u_{\text{imp}}=\sqrt{E}\,\phi^{\frac{d-1}{2}}\left(\frac{\lambda}{R}\right)^{\frac{d}{2}+1-h_{d}}$ (22)